Which of the following statements is FALSE?

An aging schedule categorizes accounts by the number of days they have been on the firm's books.

The lower the discount percentage offered, the greater the cost of forgoing the discount and using trade credit.

The collection process, although sometimes difficult, is also expensive in terms of out-of-pocket expenses.

Knowledge of the payments pattern is also useful for forecasting the firm's working capital requirements.

After a firm decides on its credit standards, it must next establish its credit terms.
Answer #1

Answer: The second statement is the false statement
Explanation:

Cost of trade credit = (1+d÷(1-d))^(365÷n)-1
Here d refers to the discount rate and n refers to the number of days after discount period

When the effective interest rate 2% discount
Cost of trade credit = (1+d÷(1-d))^(365÷n)-1
=(1+0.02/(1-0.02))^(365/25)-1
=1.020408163^(365/25)-1
=0.34307 or 34.31% (Rounded to two decimal places)


When the effective interest rate 3% discount
=(1+0.03/(1-0.03))^(365/25)-1
=(1.030927835)^(365/25)-1
=0.560 or 56.00% (Rounded to two decimal places)

So, its not true that "lower the discount percentage offered, the greater the cost of forgoing the discount and using trade credit".